When a solution of NaOH is neutralized by hydrochloric acid, the solution gets hot.

Calculate the temperature change of the solution (in K).


Mass of combined solution = 125 g

Heat change of the solution = 3490 J

Specific heat capacity of the solution = 4.18 J/g·K


Give your answer to three significant figures.


1
Expert's answer
2020-08-25T12:19:29-0400

From a formula for heat change:

Q=m×c×∆T

∆T=Q/m×c

∆T=3490/125×4.18=6.679(K).